OpenStudy (anonymous):

The Limit as t approaches 0 of 3/Square root of t.

OpenStudy (anonymous):

\[\lim_{t\to0^+}\frac{3}{\sqrt t}~~?\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

It's infinity. As \(t\) gets smaller (approaching 0), \(\sqrt t\) gets smaller, so \(\dfrac{1}{\text{small number}}\to\infty\).

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw